OpenStudy (anonymous):

I have one more question does the letter i is not also use for the symbol +- sign together

OpenStudy (anonymous):

i = sqrt (-1) sqrt(-4) = 2i 3 +/- sqrt (-4) = 3 +/- 2i

OpenStudy (anonymous):

Like for instance where you have the 3 +/- can you not replace that with an i .

OpenStudy (anonymous):

3 +/- sqrt (-4) = 3 +/- 2i is just an example cuz i know ur working on quadratic let b = -3 a = 1/2 sqrt (D) = sqrt (-4) x = 3 +/- sqrt (-4) x[1] = 3+ 2i x[2] = 3-2i u cant replace 3 +/- with i because i means sqrt (-1) sqrt (-4) = sqrt (4)*sqrt(-1) = 2*i = 2i

OpenStudy (angela210793):

i is used at complex no. only...or vectors perhaps :)